Use substitution to solve
2x - 4y = 14
x = -2y + 7

Found 2 solutions by stanbon, richwmiller:
Answer by stanbon(75887)

Use substitution to solve
2x - 4y = 14
x = -2y + 7
---------
Substitute for "x" and solve for "y"::
2(-2y+7)-4y = 14
-4y+14-4y = 14
-8y = 0
y = 0
-----
Solve for "x"::
2x -4*0 = 14
2x = 14
x = 7
============
Cheers,
Stan H.
============

Answer by richwmiller(17219)

plug -2y+7 for x in
x=7 y=0